# Additive manufacturing of a 3D-segmented plastic scintillator detector for tracking and calorimetry of elementary particles

### v1

@@ IV.2 Particle detector fabrication
The SuperCube was produced in two separate processes; the fabrication of a $5\times 5$ matrix layer of an optically reflective frame and the forming of the PS, each using a different mechanical setup. The two production steps were alternated to build a layer-by-layer three-dimensional $5\times 5\times 5$ SuperCube. The reflective frame with holes for the WLS fibers was manufactured via FDM without the use of support structure (Fig. 7 a). The PS was formed in three sub-processes. First, metal rods were placed through the holes to create circular voids for the WLS fibers to be inserted, then the square-shaped volume was rapidly filled in a bottom-to-top motion with a customized extrusion setup combined with a pressurised plate to keep the melt constrained in the cavity (Fig. 7 b) and thirdly, a heated punch was pressed on top of the PS to plane its surface such that the next matrix-layer could be fabricated on top of it (Fig. 7 c). Finally, the top of the voxel was closed by a top layer and the WLS fibers were inserted through the already manufactured holes (Fig. 7 d).
@@ IV.2.1 Reflective Frame Fabrication
A 1.75 mm polycarbonate-polytetrafluoroethylene-blend (PC-PTFE) filament from Rosa3D filaments [58] was used for the frame fabrication. This material provided the required combination of good optical reflectivity to contribute to the performance of hte detector and high thermal resistance to ensure the dimensional integrity of the geometry during the forming of the PS.
The structure was printed with a nozzle temperature of 280°C, a bed temperature of 115°C for the first matrix layer and a chamber temperature of 55°C to reduce the risk of warping. After the first filling cycle, the bed temperature was set to 75°C at all time to avoid permanently turning the scintillation material opaque, as was observed in previous studies [38]. The layer height was set to 0.2 mm, the print speed at 2000 mm/min, and cooling was only enabled during the layers that incorporated circular holes to ensure accurate round geometries without the use of support structure. In FDM, the thickness of horizontally-printed walls (parallel to the print bed) is very close to the designed value and was not further investigated. However, 20 thickness measurements of vertically-printed walls (perpendicular to the print bed) were taken using a caliper to show the accuracy and precision of this manufacturing step.
@@ IV.2.2 Plastic scintillator forming
To maximize the particle detector performance, the PS needed to be transparent and have a high fill factor throughout the voxel to increase the active material, and around the metal rods to have a large contact surface between the scintillation material and the WLS fibers for a better light capture efficiency. The fill density is dependent, besides the viscosity of the polymer, on the magnitude of the mass flow rate exiting the nozzle. When increased, the system distributes the melt within the cavity with more energy, pushing the material towards the boundaries with more pressure. This shortens the distribution time of the melt without the polymer being able to cool down and recrystallize before it has reached the outer walls of the reflective shell. An even distribution of scintillation material within the entire cavity shall be achieved by an elongated extrusion nozzle that could reach the bottom of the reflective shell. The liquid polymer was deposited in vertical bottom-to-top movement, creating a single blob, wherefrom the melt spread towards the boundaries. At the top, a spring-loaded, polished, stainless steel plate constrained the melt pool within the cavity (Fig. 8 a). The plate included a hole for the nozzle to move freely up and down, vent holes for air to escape during voxel-filling and a concave dome for a temporarily overfill that compensated for an ensuing material shrinkage (Fig. 8 b).
Figure 8: a) Depiction of the extrusion system components modified for the filling of scintillation material into the reflective cavity. b) Geometry of the pressurized plate that is pressed onto the cavity tho keep the melt pool restrained in fill volume.
A CFD analysis using ANSYS Fluent was performed to determine material requirements of the melting components: material of heat block and extrusion nozzle, geometry and material of the heat break and the feeding tube of the filament; and the process parameters heat block temperature and extrusion speed (Fig.9 a).
These process elements were balanced to attain a PS outflow temperature at the nozzle orifice of around 230°C to preserve the scintillation properties established in [36, 38]; to reach the maximum possible material throughput velocity in order to obtain a high fill density of the PS volume by more forcefully spreading the melted polymer throughout the reflective cavity before it solidifies, which is achieved by elevating the heat block temperature (above the usual working temperature) such that the PS is melted to its core at the exit point; and to simultaneously handle the increased heat creep from the high temperature heat block towards the whole extrusion system, which could lead to clogging followed by the failure of the entire injection process.
@@ IV.3 Detector setup and analysis method for cosmic data
The $5\times 5\times 5$ SuperCube was instrumented with fifty readout channels, twenty-five on each view, as shown in Fig. 4. Kuraray Y11 double-cladding 1 mm diameter WLS fibers [40] along the orthogonal X and Y directions allow to uniquely identify the cubes crossed by a charged particle. WLS fibers absorb the blue light produced by the PS cube and isotropically emit photons in the green wavelength inside their core. The different refractive index of the outer cladding geometrically traps and guides the photons towards the SiPMs, thanks to an attenuation length of more than 4 meters. Hamamatsu Multi-Pixel Photon Counters (MPPCs) S13360-1325CS [59] were coupled with individual WLS fibers, on one of the two ends. The operating voltage was set to the one recommended by the supplier for each MPPC. The coupling was ensured by black 3D printed optical connectors and enhanced by fixing the WLS fiber with EJ-500 optical glue [60] and pushing the MPPC toward the polished fiber end with a soft black foam acting like a spring. With a nominal photodetection efficiency of 25%, the analogue charge signal of each MPPC is proportional, to the first order, to the number of photons produced in the PS cubes crossed by the corresponding WLS fiber. The charge analogue signal of each MPPC is collected with micro-coaxial cables and read out by one CAEN FERS DT5202 front-end (FE) board [61], used to digitize the signal. The charge value corresponding to the highest signal peak of each independent channel is measured and converted to analogue-to-digital converter (ADC) units. Then, from the measured number of ADC units, the number of photoelectrons (PE), i.e., the primary electrons originated via photoelectric effect by the visible photons impinging on the MPPC active area, is computed. Such conversion was obtained with a large data sample collected by exposing the prototype to a $\text{Sr}^{90}$ radioactive source. The data could show a clear multi-peak structure, each one corresponding to a different number of PEs. The MPPC gain could be extracted as the distance between adjacent peaks in units of ADC per PE with a multi-Gaussian distribution fit and the position of the first three peaks could be obtained. Then, the peak positions were fitted with a linear function to extract the gain and pedestal. The above explained procedure was repeated for every channel. The different gains were found to be relatively uniform around 50 ADC/PE. Examples of detected cosmic particles can be found in Fig. 4.
@@
The performance of the SuperCube produced with FIM was compared with a reference 2D matrix of cubes produced with injection molding. It consisted of a single layer of $5\times 5$ optically-isolated PS $1~{}\text{cm}^{3}$ cubes, thus a geometry analogous to the SuperCube but two dimensional. In the same way as for the SuperCube, both orthogonal views of the reference sample were read out with Kuraray Y11 WLS fibers, coupled with the same type of MPPC following the same procedure as described above. A more detailed description of the reference prototype as well as of its particle detection performance can be found in [30]. The reference prototype was placed on top of the SuperCube and, read out with the same CAEN FERS DT5202 board, was used to trigger the cosmic particles and compare the light yield and crosstalk with the one of the SuperCube.
@@
Both the FIM and the injection molding prototypes were characterised by analyzing cosmic particle events and the respective light yield and crosstalk were compared. A total of about 63 hours cosmic data was collected. Typical cosmic events are minimum ionising particles, crossing the prototype every few seconds with an expected energy loss in PS of about 1.8 MeV/cm. Given their angular distribution, approximately equal to $\cos^{2}\theta_{\text{azim}}$ where $\theta_{\text{azim}}$ is the azimuth angle, the most common signature consists to the one of a vertical MIP leaving a track that starts from the top layer (the reference prototype) and ends at the bottom layer of the SuperCube, as shown in Fig. 4. The track 3D hits were reconstructed layer by layer, by determining the XY coordinate with the maximum light yield channel in both views. In each electronic channel a threshold of 500 ADC, i.e. about 10 PE, was applied. After the track reconstruction, a data set containing only vertical tracks, i.e. crossing both the reference prototype and the SuperCube in the same column of cubes, was selected to have a particle path length of about 1 cm in each PS cube, hence a comparable energy loss. In Fig. 5, a scintillation light yield of about 29 PE per cube is shown for both the SuperCube and the reference sample. The cube-to-cube light crosstalk was derived from the light yield ratio between neighboring channels within the same layer perpendicular to the track direction. As shown in Fig. 5, a slightly higher crosstalk was measured for the SuperCube . However, it remained constrained at the few % level, thus acceptable for particle tracking and calorimetry measurements.

### v2

@@ IV.2 Particle detector fabrication
The SuperCube was produced in two separate processes; the fabrication of a $5\times 5$ matrix layer of an optically reflective frame and the forming of the PS, each using a different mechanical setup. The two production steps were alternated to build a layer-by-layer three-dimensional $5\times 5\times 5$ SuperCube. The reflective frame with holes for the WLS fibers was manufactured via FDM without the use of support structure (Fig. 7 a). The PS was formed in three sub-processes. First, metal rods were placed through the holes to create circular voids for the WLS fibers to be inserted, then the square-shaped volume was rapidly filled in a bottom-to-top motion with a customized extrusion setup combined with a pressurized plate to keep the melt constrained in the cavity (Fig. 7 b) and thirdly, a heated punch was pressed on top of the PS to plane its surface such that the next matrix-layer could be fabricated on top of it (Fig. 7 c). Finally, the top of the voxel was closed by a top layer and the WLS fibers were inserted through the already manufactured holes (Fig. 7 d).
@@ IV.2.1 Reflective Frame Fabrication
A 1.75 mm polycarbonate-polytetrafluoroethylene-blend (PC-PTFE) filament from Rosa3D filaments [79] was used for the frame fabrication. This material provided the required combination of good optical reflectivity to contribute to the performance of the detector and high thermal resistance to ensure the dimensional integrity of the geometry during the forming of the PS.
The structure was printed with a nozzle temperature of 280°C, a bed temperature of 115°C for the first matrix layer and a chamber temperature of 55°C to reduce the risk of warping. After the first filling cycle, the bed temperature was set to 75°C at all time to avoid permanently turning the scintillation material opaque, as was observed in previous studies [55]. The reflective structure was fabricated using a 0.4 mm diameter nozzle with a fill factor of 100% for both horizontal and vertical walls. The layer height was set to 0.2 mm, the print speed to 2000 mm/min, and cooling was activated only during the layers incorporating circular holes to ensure accurate round geometries without the need for support structures. In FDM, the thickness of horizontally-printed walls (parallel to the print bed) is very close to the designed value and was not further investigated. However, 20 thickness measurements of vertically-printed walls (perpendicular to the print bed) were taken using a caliper to show the accuracy and precision of this manufacturing step.
The measured transmission for bottom, top and side walls of the white reflective shell, taken in air, is shown in Fig. 8. Data were collected with a monochromatic light source and read out with an integrated sphere and a photodiode (Hamamatsu S1337-1010BQ).
Figure 8: Measured transmittance of white reflector sheets with the same thickness but different print orientation compared to the printing bed. Vertical (side) walls in red and horizontal (top and bottom) walls in blue of a PS voxel. The experimental data were collected in air.
@@ IV.2.2 Plastic scintillator forming
To optimize the performance of the particle detector, the material required high transparency and a consistent fill factor throughout the voxel to maximize the active material volume. Additionally, it was crucial for the PS to surround the metal rods, ensuring a substantial contact area between the scintillation material and the WLS fibers to enhance light capture efficiency. The fill density was influenced not only by the viscosity of the polymer but also by the magnitude of the mass flow rate exiting the nozzle. Increasing the flow rate resulted in a more energetic distribution of the melt within the cavity, exerting greater pressure on the material towards the boundaries. This accelerated the distribution of the melt without allowing the polymer to cool and recrystallize before reaching the outer walls of the reflective shell. An elongated extrusion nozzle, capable of reaching the bottom of the reflective shell, facilitated an even distribution of scintillation material throughout the entire cavity. The liquid polymer was deposited through the 1.8 mm wide nozzle orifice in a vertical bottom-to-top motion, forming a single blob from which the melt spread towards the boundaries. Positioned at the top, a spring-loaded, polished stainless steel plate constrained the melt pool within the cavity. The plate incorporated a hole to enable unrestricted vertical movement of the nozzle, along with vent holes designed to release air during voxel-filling. Additionally, it featured a concave dome to provide temporary overfill, effectively compensating for any subsequent material shrinkage.
Figure 9: a) Depiction of the extrusion system components modified for the filling of scintillation material into the reflective cavity. b) Geometry of the pressurized plate that is pressed onto the cavity to keep the melt pool restrained in fill volume.
A CFD analysis using ANSYS Fluent was performed to determine the material requirements of the melting components heat block and extrusion nozzle, the geometry and material of the heat break and the feeding tube of the filament, and the process parameters heat block temperature and extrusion speed (Fig.10 a).
These process elements were balanced to attain a PS outflow temperature at the nozzle orifice of around 230°C to preserve the scintillation properties established in [53, 55] and to reach the maximum possible material throughput velocity in order to obtain a high fill density of the PS volume by more forcefully spreading the melted polymer throughout the reflective cavity before it solidifies. This is achieved by elevating the heat block temperature (above the usual working temperature) such that the PS is melted to its core at the exit point. Simultaneously, the increased heat creep from the high-temperature heat block towards the whole extrusion system needs to be controlled, such that system clogging followed by the failure of the entire injection process is prevented.
@@ IV.3 Detector setup and analysis method for cosmic data
The $5\times 5\times 5$ SuperCube was instrumented with fifty readout channels, twenty-five on each view, as shown in Fig. 4. Kuraray Y11 double-cladding 1 mm diameter WLS fibers [58] along the orthogonal X and Y directions allow for unique identification of the cubes crossed by a charged particle. WLS fibers absorb the blue light produced by the PS cube and isotropically emit photons in the green wavelength inside their core. The different refractive index of the outer cladding geometrically traps and guides the photons towards the SiPMs, thanks to an attenuation length of more than 4 meters. Hamamatsu Multi-Pixel Photon Counters (MPPCs) S13360-1325CS [80] were coupled with individual WLS fibers, on one of the two ends. The operating voltage was set to the one recommended by the supplier for each MPPC. The coupling was ensured by black 3D printed optical connectors and enhanced by fixing the WLS fiber with EJ-500 optical glue [81] and pushing the MPPC toward the polished fiber end with a soft black foam acting like a spring. The other end of the fiber was cut at $90^{\circ}$ and polished. With a nominal photodetection efficiency of 25%, the analogue charge signal of each MPPC is proportional, to the first order, to the number of photons produced in the PS cubes crossed by the corresponding WLS fiber. The charge analogue signal of each MPPC is collected with micro-coaxial cables and read out by one CAEN FERS DT5202 front-end (FE) board [82], used to digitize the signal. The charge value corresponding to the highest signal peak of each independent channel is measured and converted to analogue-to-digital converter (ADC) units. Then, from the measured number of ADC units, the number of photoelectrons (PE), i.e., the primary electrons originated via photoelectric effect by the visible photons impinging on the MPPC active area, is computed. Such conversion was obtained with a large data sample collected by exposing the prototype to a ${}^{90}\text{Sr}$ radioactive source. The data could show a clear multi-peak structure, each one corresponding to a different number of PEs. The MPPC gain could be extracted as the distance between adjacent peaks in units of ADC per PE with a multi-Gaussian distribution fit and the position of the first three peaks could be obtained. Then, the peak positions were fitted with a linear function to extract the gain and pedestal. Since the recommended working bias voltage provided by the supplier was applied to each individual channel of the MPPC, the gain distribution was observed to be relatively uniform, averaging around 50 ADC/PE with a standard deviation of 1.21 ADC/PE. This procedure was iterated for each channel, employing its specific gain and pedestal value to convert the measured light yield into units of PE. Examples of detected cosmic particles can be found in Fig. 4.
@@
The performance of the SuperCube produced with FIM was compared with a reference 2D matrix of cubes produced with cast polymerization. It consisted of a single layer of $5\times 5$ optically-isolated PS $1~{}\text{cm}^{3}$ cubes, thus a geometry analogous to the SuperCube but two dimensional. In the same way as for the SuperCube, both orthogonal views of the reference sample were read out with Kuraray Y11 WLS fibers, coupled with the same type of MPPC following the same procedure as described above. A more detailed description of the reference prototype as well as of its particle detection performance can be found in [41]. The reference prototype was placed on top of the SuperCube and, read out with the same CAEN FERS DT5202 board, was used to trigger the cosmic particles and compare the light yield and crosstalk with the one of the SuperCube.
@@
Both the FIM and the cast polymerization prototypes were characterized by analyzing cosmic particle events and the respective light yield and crosstalk were compared. A total of about 63 hours cosmic data was collected. Typical cosmic events are minimum ionising particles, crossing the prototype every few seconds with an expected energy loss in PS of about 1.8 MeV/cm. Given their angular distribution, approximately equal to $\cos^{2}\theta_{\text{azim}}$ where $\theta_{\text{azim}}$ is the azimuth angle, the most common signature consists to the one of a vertical MIP leaving a track that starts from the top layer (the reference prototype) and ends at the bottom layer of the SuperCube, as shown in Fig. 4. The track 3D hits were reconstructed layer by layer, by determining the XY coordinate with the maximum light yield channel in both views. In each electronic channel a threshold of 500 ADC, i.e., about 10 PE, was applied. After the track reconstruction, a data set containing only vertical tracks was selected to have a particle path length of about 1 cm in each PS cube, hence a comparable energy loss. These tracks cross both the reference prototype and the SuperCube in the same column of cubes. In Fig. 5, a scintillation light yield of about 29 PE per cube is shown for both the SuperCube and the reference sample. The cube-to-cube light crosstalk was derived from the light yield ratio between neighboring channels within the same layer perpendicular to the track direction. As shown in Fig. 5, a slightly higher crosstalk was measured for the SuperCube. However, it remained constrained at the few % level, thus acceptable for particle tracking and calorimetry measurements.